1) 2) 3)

1)

1) 2) 3)

2)

3)
Answer #1


Speed vx = dx/dt = (1.5*t^2 - 4t)

vy =dy/dt = t-2

at t = 4.5S

vx = (1.5*4.5*4.5)-(4*4.5) = 12.375 m/s

vy = 4.5-2 = 2.5 m/s

then speed is v = sqrt(vx^2+vy^2) = sqrt(12.375^2+2.5^2) = 12.625 m/s

----------------------------------------------------

pART d)


theta = atan(vy/vx) = atan(2.5/12.375) = 11.2 degrees

----------------------------------------------

2) A) velocity at the top of the hole is v = 18 m/s
rock's velocity at the bottom of the hole is v = sqrt(u^2+(2*g*h)) = sqrt(18^2+(2*9.8*10)) = 22.8 m/s

B) T = (2*u/g)+((v-u)/g) = (2*18/9.81)+((22.8-12.625) = 13.8 S


3) partA)

ax = d^2x/dt^2 = 8 m/s

ay = -6m/s

magitude of acceration is sqrt(ax^2+ay^2) = sqrt(8^2 + 22.8 m/s) = 1459.2 m/s^2


direction is thete =atam(vy/ve) then a = atna(vy/vx) = atan(0.5)= 26.56 V

part C) v at 2 S

vx = 4 m/s
and Vb = -2 m/s

magnitude is sqrt4^2+2^2) = 2*sqrt(5)


theta = atan(vy/v e) = atan(-2/4 ) = 26.56 degrees